OpenStudy (just_one_last_goodbye):

Which ratio forms a proportion with 30/36?

OpenStudy (just_one_last_goodbye):

3/5 5/6 15/12 6/5

OpenStudy (anonymous):

b

OpenStudy (tylerd):

30/36 = 15/18 = 5/6 notice the pattern?

OpenStudy (tylerd):

half of 30 is 15, half of 36 is 18, 1/3 of 15 is 5 and 1/3 of 18 is 6

OpenStudy (idealist10):

5/6

OpenStudy (idealist10):

30/36=5/6

OpenStudy (just_one_last_goodbye):

thanks Guys ^_^ @TylerD i would have given u a medal but i felt bad that idk didn't get any and u got 2